# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ Abstract
Stochastic gradient descent (SGD) has become a cornerstone of neural network optimization, yet the noise introduced by SGD is often assumed to be uncorrelated over time, despite the ubiquity of epoch-based training. In this work, we challenge this assumption and investigate the effects of epoch-based noise correlations on the stationary distribution of discrete-time SGD with momentum, limited to a quadratic loss. Our main contributions are twofold: first, we calculate the exact autocorrelation of the noise for training in epochs under the assumption that the noise is independent of small fluctuations in the weight vector, and find that SGD noise is anti-correlated in time. Second, we explore the influence of these anti-correlations on SGD dynamics. We find that for directions with a curvature greater than a hyperparameter-dependent crossover value, the results for uncorrelated noise are recovered. However, for relatively flat directions, the weight variance is significantly reduced, and our variance prediction leads to a considerable reduction in loss fluctuations as compared to the constant weight variance assumption.
@@ 1 Introduction
Initially developed to address the challenges of computational efficiency in neural networks, stochastic gradient descent (SGD) has exhibited exceptional effectiveness in managing large datasets compared to the full-gradient methods [1]. It has since garnered widespread acclaim in the machine learning domain [2], with applications spanning image recognition [3, 4, 5], natural language processing [6, 7], and mastering complex games beyond human capabilities [8]. Alongside its numerous variants [9, 10, 11], SGD remains the cornerstone of neural network optimization.
@@
To deepen our understanding, multiple studies have investigated the limiting dynamics of neural network weights during the latter stages of training [27, 28]. Of particular interest is the behavior of weight fluctuations in proximity to a minimum of the loss function [29, 15, 30]. Several authors found empirical evidence that the covariance matrix $\bf C$ of SGD is proportional to the Hessian matrix $\bf H$ of the loss function [20, 21, 22, 23, 31, 18]. Consequently, theory posits that the stationary covariance matrix of weights $\bf\Sigma$ exhibits isotropy for sufficiently small learning rates [15, 28, 30]. Nevertheless, a recent empirical investigation [32] identified profound anisotropy in $\bf\Sigma$.
In this work, we delve into both theoretical and empirical analyses of weight fluctuations during the later stages of training, accounting for the emergence of anti-correlations in the noise produced by SGD, which stem from the prevalent epoch-based learning schedule. As a result of these anti-correlations, we discover that the covariance matrix $\bf\Sigma$ displays anisotropy in a subspace of weight directions corresponding to Hessian eigenvectors with small eigenvalues, while maintaining the isotropy of $\bf\Sigma$ in directions associated with Hessian eigenvectors possessing large eigenvalues.
@@
We make the following contributions:
  • We uncover generic anti-correlations in SGD noise, resulting from the common practice of drawing examples without replacement. These anti-correlations lead to a smaller-than-anticipated weight variance in some directions.
@@
  • We calculate the autocorrelation function of SGD noise, assuming that the noise is independent of small fluctuations in the weight vector over time. Given that the noise sums to zero over an epoch where each example is shown once, we find anti-correlations of SGD noise.
@@
  • Building on the computed autocorrelation function, we develop a theory that elucidates the relationships between variances for different Hessian eigenvectors. The weight space can be partitioned into two groups of Hessian eigenvectors, with the well-known isotropic variance prediction holding for eigenvectors whose Hessian eigenvalues exceed a hyperparameter-dependent crossover value $\lambda_{\textrm{cross}}$. For eigenvectors with smaller eigenvalues, the variance is proportional to the Hessian eigenvalue, smaller than the isotropic value. This theory accounts for SGD’s discrete nature, avoiding continuous-time approximations and incorporating heavy-ball momentum.
@@
  • For a given Hessian eigenvector, an intrinsic correlation time of update steps exists, introduced by the loss landscape and proportional to $1/\lambda_{i}$, where $\lambda_{i}$ represents the corresponding Hessian eigenvalue. Additionally, a fixed noise correlation time $\tau_{\textrm{SGD}}$ is present, on the order of one epoch and direction-independent. The smaller of these two timescales governs the dynamic behavior for a given Hessian eigenvector, determining the actual correlation time $\tau_{i}$ of the update steps. As weight variances are proportional to $\tau_{i}$, the two distinct variance-curvature relationships arise.
@@
  • Our theoretical predictions are validated through the analysis of a neural network’s training within a subspace of its top Hessian eigenvectors. Utilizing Hessian eigenvectors proves advantageous over the principal component analysis of the weight trajectory used in a prior empirical study [32], where finite size effects induced significant artifacts.
  • We investigate the effect of weight variance on training error in neural networks, focusing on Hessian eigenvectors and weight fluctuations. Our analysis, based on a quadratic loss model with Gaussian weight fluctuations, demonstrates that the reduction of weight fluctuations due to noise anti-correlations leads to a significant reduction of loss fluctuations. In a case study with the LeNet architecture, analyzing the 5,000 largest Hessian eigenvalues, we observe a 62% reduction in loss fluctuation compared to models assuming constant weight variance. This highlights the importance of considering a broader range of Hessian eigendirections in understanding and optimizing neural network training dynamics.
@@ 2 Background
We consider a neural network characterized by its weight vector, $\bm{\mathbf{\theta}}\in\mathbb{R}^{d}$. The network is trained on a set of $N$ training examples, each denoted by $x_{n}$, where $n$ ranges from 1 to $N$. The loss function, defined as ${L(\bm{\mathbf{\theta}})\coloneqq\frac{1}{N}\sum_{i=1}^{N}l(\bm{\mathbf{\theta}},x_{n})}$, represents the average of individual losses incurred for each training example $l(\bm{\mathbf{\theta}},x_{n})$.
@@
$\bm{\mathbf{g}}_{k}(\bm{\mathbf{\theta}})$ $=\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\bm{\mathbf{\nabla}}l(\bm{\mathbf{\theta}},x_{n})\ ,$ (1a)
$\bm{\mathbf{v}}_{k}$ $=-\eta\bm{\mathbf{g}}_{k}(\bm{\mathbf{\theta}}_{k-1})+\beta\bm{\mathbf{v}}_{k-1}\ ,$ (1b)
$\bm{\mathbf{\theta}}_{k}$ $=\bm{\mathbf{\theta}}_{k-1}+\bm{\mathbf{v}}_{k}\ .$ (1c)
Here, $k$ signifies the discrete update step index, $\eta$ is the learning rate, and $\beta$ is the momentum parameter. The stochastic gradient at each step is computed with respect to a batch of $S\ll N$ random examples. Each batch is denoted by $\mathcal{B}_{k}=\{n_{1},...,n_{S}\}$, where $n_{j}\in\{1,\dots,N\}$. The training process is structured into epochs. During each epoch, every training example is used exactly once, implying that the examples are drawn without replacement and do not recur within the same epoch.
In the realm of SGD as opposed to full gradient descent, we introduce noise, denoted as $\vec{\delta g}_{k}(\vec{\theta})\coloneqq\vec{g}_{k}(\vec{\theta})-\vec{\nabla}L(\vec{\theta})$, with a covariance matrix $\bm{\mathbf{C}}(\vec{\theta})\coloneqq\textrm{cov}\bigl{(}\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k}(\vec{\theta})\bigr{)}$. The noise covariance matrix is known to be proportional to the gradient sample covariance matrix $\vec{C_{0}}(\vec{\theta})\coloneqq\frac{1}{N-1}\sum_{n=1}^{N}\vec{\nabla}\left[l(\bm{\mathbf{\theta}},x_{n})-L(\vec{\theta})\right]\vec{\nabla}^{\top}\left[l(\bm{\mathbf{\theta}},x_{n})-L(\vec{\theta})\right]$. A detailed derivation of this relation, particularly for the case of drawing without replacement, is provided in Appendix E.
Our primary focus lies on the asymptotic or limiting covariance matrix of the weights $\bm{\mathbf{\Sigma}}\coloneqq\textrm{cov}\bigl{(}\vec{\theta}_{k},\vec{\theta}_{k}\bigr{)}$. That means we are interested in the covariance computed over an infinite run of SGD optimization, and not the covariance for a fixed update step $k$ computed over multiple runs of SGD optimization (see Appendix A).
To better understand and explain the behavior of the weight variances, we further examine the covariance matrix of the velocities $\bm{\mathbf{\Sigma_{v}}}\coloneqq\textrm{cov}\bigl{(}\vec{v}_{k},\vec{v}_{k}\bigr{)}$, hypothesize that these two matrices commute, and proceed to explore their ratio $\bm{\mathbf{\Sigma}}/\bm{\mathbf{\Sigma_{v}}}$. The eigenvalues of this matrix ratio are denoted as $\tau_{i}$ because, under general assumptions, they equate to the velocity correlation time of the corresponding eigendirection (see Section 4.2).
@@ 3 Related Work
3.1 Hessian Noise Approximation
The equivalence between the gradient sample covariance $\vec{C_{0}}$ and the Hessian matrix of the loss $\vec{H}$ is an approximation that frequently appears in the literature [15, 22, 30]. Numerous theoretical arguments have indicated that when the output of a neural network closely matches the example labels, these two matrices should be similar [33, 15, 20, 21, 23]. However, even a slight deviation between network predictions and labels can theoretically disrupt this relationship, as highlighted by Thomas et al. [31]. Despite this, empirical observations suggest a strong alignment between the gradient sample covariance and the Hessian matrix near a minimum.
Zhang et al. [22] pursued this line of thought and evaluated the assumption numerically. They examined both matrices in the context of a particular basis that presents both high and low curvature across different directions. Their findings indicated a close match between curvature and gradient variance in a given direction for a convolutional image recognition network, and a reasonably good relationship for a transformer model.
Meanwhile, Thomas et al. [31] presented both a theoretical argument and empirical evidence across different architectures for image recognition. Although they did not discover an exact match between the two matrices, they did observe a proportionality between them, indicated by a high cosine similarity.
Xie et al. [18] also conducted an investigation with an image recognition network. In the eigenspace of the Hessian matrix, they plotted entries within a specific interval against corresponding entries from the gradient sample covariance and found a close match.
For our theoretical considerations, it is crucial to assume that both matrices commute, $\left[\bm{\mathbf{C_{0}}},\bm{\mathbf{H}}\right]=0$. Additionally, we hypothesize that $\bm{\mathbf{C_{0}}}\propto\bm{\mathbf{H}}$, to derive power law predictions for the weight variance.
@@ 3.2 Limiting Dynamics and Weight Fluctuations
Various studies have scrutinized the limiting dynamics, often modeling SGD as a stochastic differential equation (SDE). Commonly, researchers such as Mandt et al. [29] and Jastrzębski et al. [15] approximate the loss near a minimum as a quadratic function and present the SDE as a multivariate Ornstein-Uhlenbeck (OU) process. This process proposes a stationary weight distribution with Gaussian weight fluctuations. Jastrzębski et al. [15] further assume the Hessian noise approximation, observing under these conditions that the weight fluctuations are isotropic. Kunin et al. [28] who also incorporate momentum into their analysis, predict and empirically verify isotropic weight fluctuations. Chaudhari & Soatto [34] also investigate the SDE but without the assumption of a quadratic loss, nor that it reached equilibrium. They gain insights via the Fokker-Plank equation.
Alternatively, some studies derive relations from a stationarity assumption instead of a continuous time approximation [27, 30, 25]. Yaida [27] assumes that the weight trajectory follows a stationary distribution and derives general fluctuation-dissipation relations from that. Liu et al. [30] go further to assume a quadratic loss function, leading them to derive exact relations for the weight variance of SGD with momentum. If the additional Hessian noise approximation is made, their results also predict the weight variance to be approximately isotropic, except in directions where the product of learning rate and Hessian eigenvalue is significantly high.
Such computed weight variances are explicitly applied in various contexts, such as computing the escape rate from a minimum or assessing the approximation error in SGD, which captures the additional training error attributed to noise [30].
Feng & Tu [32] present a phenomenological theory based on their empirical findings, which also account for flat directions, unlike Kunin et al. [28]. They describe a general inverse variance-flatness relation, analyzing the weight trajectory of different image recognition networks via principal component analysis. They discovered a power law relationship between the curvature of the loss and the weight variance $\sigma_{\theta,i}^{2}$ in any given direction, where a higher curvature corresponds to a higher variance. They also observed that both the velocity variance $\sigma_{v,i}^{2}$ and the correlation time $\tau_{i}$ are larger for higher curvatures.
In our approach, we do not make the continuous time approximation but base our results on the assumption that the weights adhere to a stationary distribution near a quadratic minimum.

### v3

# Anti-Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances in Flat Directions

@@ Abstract
Stochastic Gradient Descent (SGD) has become a cornerstone of neural network optimization due to its computational efficiency and generalization capabilities. However, the gradient noise introduced by SGD is often assumed to be uncorrelated over time, despite the common practice of epoch-based training where data is sampled without replacement. In this work, we challenge this assumption and investigate the effects of epoch-based noise correlations on the stationary distribution of discrete-time SGD with momentum. Our main contributions are twofold: First, we calculate the exact autocorrelation of the noise during epoch-based training under the assumption that the noise is independent of small fluctuations in the weight vector, revealing that SGD noise is inherently anti-correlated over time. Second, we explore the influence of these anti-correlations on the variance of weight fluctuations. We find that for directions with curvature of the loss greater than a hyperparameter-dependent crossover value, the conventional predictions of isotropic weight variance under stationarity, based on uncorrelated and curvature-proportional noise, are recovered. Anti-correlations have negligible effect here. However, for relatively flat directions, the weight variance is significantly reduced, leading to a considerable decrease in loss fluctuations compared to the constant weight variance assumption. Furthermore, we present a numerical experiment where training with these anti-correlations enhances test performance, suggesting that the inherent noise structure induced by epoch-based training may play a role in finding flatter minima that generalize better.
  • 8 October 2025
Keywords: Stochastic Gradient Descent, Asymptotic Analysis, Discrete Time, Hessian
@@ 1 Introduction
Initially developed to address the challenges of computational efficiency in neural networks, stochastic gradient descent (SGD) has exhibited exceptional effectiveness in managing large datasets compared to the full gradient methods [1]. It has since garnered widespread acclaim in the machine learning domain [2], with applications spanning image recognition [3, 4, 5], natural language processing [6, 7], and mastering complex games beyond human capabilities [8]. Alongside its numerous variants [9, 10, 11], SGD remains the cornerstone of neural network optimization.
@@
To deepen our understanding of neural network training dynamics, several studies have investigated the limiting behavior of network weights during the later stages of training [27, 28]. Of particular interest is the behavior of weight fluctuations near a minimum of the training loss function [29, 15, 30]. The variance of these fluctuations reveals how strongly different directions in parameter space are explored and has been used, for example, to quantify escaping efficiency from different minima [23, 30], offering valuable insights into optimization behavior. Empirical evidence suggests that the covariance matrix $\mathbf{C}$ associated with SGD noise is proportional to the Hessian matrix $\mathbf{H}$ of the training loss function [20, 21, 22, 23, 31, 17], although this proportionality may not hold under certain conditions [25]. Here, $\mathbf{C}$ is defined as the covariance of the difference between the minibatch gradient and the full-batch gradient. Consequently, theoretical analyses predict that the stationary covariance matrix of the weights, $\mathbf{\Sigma}$, becomes isotropic when the learning rate is sufficiently small [15, 30, 28]. However, recent empirical studies [32] have identified significant anisotropy in $\mathbf{\Sigma}$.
In this work, we present both theoretical and empirical analyses of weight fluctuations during the later stages of training, accounting for the emergence of anti-correlations in the noise produced by SGD, which stem from the prevalent epoch-based learning schedule. As a result of these anti-correlations, we discover that the covariance matrix $\bf\Sigma$ displays anisotropy and is smaller than expected in a subspace of weight directions corresponding to Hessian eigenvectors with small eigenvalues, while maintaining the isotropy of $\bf\Sigma$ in directions associated with Hessian eigenvectors possessing large eigenvalues. Our theoretical predictions are validated through the analysis of a neural network’s training within a subspace of its top Hessian eigenvectors.
In addition, we demonstrate that for a small convolutional network trained on CIFAR10 the anti-correlations in SGD noise described above significantly increase the test accuracy. By linking this result to a previous study on artificially added anti-correlated noise and its benefits [33], we argue that the anti-correlations in SGD noise suppress diffusion in flat directions, and in this way may contribute to finding flatter minima with better test accuracy.
@@
  • We uncover generic anti-correlations in SGD noise that result from the common practice of drawing training examples without replacement. We calculate the autocorrelation function of SGD noise under the assumption that the noise is independent of small fluctuations in the weight vector over time. The anti-correlations arise because the noise sums to zero over an epoch where each example is presented once.
@@
  • Using the computed autocorrelation function, we develop a theory that elucidates the relationship between variances along different Hessian eigenvectors. The weight space partitions into two groups based on Hessian eigenvalues: for eigenvectors with eigenvalues exceeding a hyperparameter-dependent crossover value $\lambda_{\textrm{cross}}$, the isotropic variance prediction holds; for those with smaller eigenvalues, anti-correlations reduce the variance, resulting in values proportional to the eigenvalues and smaller than the isotropic prediction. This theory accounts for the discrete nature of SGD, avoids continuous-time approximations, and incorporates heavy-ball momentum.
@@
  • For each Hessian eigenvector, there exists an intrinsic correlation time of update steps, introduced by the loss landscape and proportional to $1/\lambda_{i}$, where $\lambda_{i}$ is the corresponding Hessian eigenvalue. Additionally, the noise anti-correlations give rise to a direction-independent noise correlation time $\tau_{\textrm{SGD}}$ on the order of one epoch. The smaller of these two timescales determines the actual correlation time $\tau_{i}$ for the update steps along a given eigenvector. Since weight variances are proportional to $\tau_{i}$, this leads to the emergence of two distinct variance-curvature relationships.
@@
  • We validate our theoretical predictions by analyzing the training of a neural network within a subspace spanned by its top Hessian eigenvectors. Utilizing Hessian eigenvectors offers advantages over principal component analysis of the weight trajectory, as used in a previous empirical study [32], where finite-size effects introduced significant artifacts.
  • We observe that drawing examples without replacement during training leads to higher test accuracy compared to drawing with replacement. By linking our findings to a prior study [33] on artificially added anti-correlated noise, we propose that the higher test accuracy is connected to the anti-correlated noise inherent in the without-replacement sampling, which may encourage convergence to flatter minima. This interpretation is further supported by our analysis of escaping efficiency, where we show that reduced variance in flat directions due to noise anti-correlations potentially stabilizes optimization in low-curvature regions and enhances generalization. In particular, for the distribution of the 5,000 largest Hessian eigenvalues derived from our LeNet case study, anti-correlations affect the escaping efficiency significantly. We observe a 62% reduction in escaping efficiency for this particular minimum with many flat directions, compared to a setup without anti-correlated noise.
@@ 2 Background
We consider a neural network characterized by its weight vector, $\boldsymbol{\mathbf{\theta}}\in\mathbb{R}^{d}$. The network is trained on a set of $N$ training examples, each denoted by $x_{n}$, where $n$ ranges from 1 to $N$. The training loss function, defined as $L(\boldsymbol{\mathbf{\theta}})\coloneqq\frac{1}{N}\sum_{n=1}^{N}l(\boldsymbol{\mathbf{\theta}},x_{n})$}, represents the average of individual losses incurred for each training example $l(\boldsymbol{\mathbf{\theta}},x_{n})$.
@@
$\boldsymbol{\mathbf{g}}_{k}(\boldsymbol{\mathbf{\theta}})=\frac{1}{S}\sum_{n\in\mathcal{B}_{k}}\boldsymbol{\mathbf{\nabla}}l(\boldsymbol{\mathbf{\theta}},x_{n})\ ,\qquad\boldsymbol{\mathbf{v}}_{k}=-\eta\boldsymbol{\mathbf{g}}_{k}(\boldsymbol{\mathbf{\theta}}_{k-1})+\beta\boldsymbol{\mathbf{v}}_{k-1}\ ,\qquad\boldsymbol{\mathbf{\theta}}_{k}=\boldsymbol{\mathbf{\theta}}_{k-1}+\boldsymbol{\mathbf{v}}_{k}\ .$ (1)
Here, $k$ signifies the discrete update step index, $\eta$ is the learning rate, and $\beta$ is the momentum parameter. The stochastic gradient at each step is computed with respect to a batch of $S\ll N$ random examples. Each batch is denoted by $\mathcal{B}_{k}=\{n_{1},\dots,n_{S}\}$, where $n_{j}\in\{1,\dots,N\}$. The training process is structured into epochs. During each epoch, every training example is used exactly once, implying that the examples are drawn without replacement and do not recur within the same epoch.
In the realm of SGD as opposed to full gradient descent, we introduce noise, denoted as $\vec{\delta g}_{k}(\vec{\theta})\coloneqq\vec{g}_{k}(\vec{\theta})-\vec{\nabla}L(\vec{\theta})$, with a covariance matrix $\boldsymbol{\mathbf{C}}(\vec{\theta})\coloneqq\textrm{cov}\bigl(\vec{\delta g}_{k}(\vec{\theta}),\vec{\delta g}_{k}(\vec{\theta})\bigr)$. The noise covariance matrix is known to be proportional to the gradient sample covariance matrix $\vec{C_{0}}(\vec{\theta})\coloneqq\frac{1}{N-1}\sum_{n=1}^{N}\vec{\nabla}\left[l(\boldsymbol{\mathbf{\theta}},x_{n})-L(\vec{\theta})\right]\vec{\nabla}^{\top}\left[l(\boldsymbol{\mathbf{\theta}},x_{n})-L(\vec{\theta})\right]$. A detailed derivation of this relation, particularly for the case of drawing without replacement, is provided in ˜D.
As in previous studies, our primary theoretical focus is on the asymptotic – or limiting – covariance matrix of the weights, denoted by $\boldsymbol{\mathbf{\Sigma}}\coloneqq\textrm{cov}\bigl(\vec{\theta}_{k},\vec{\theta}_{k}\bigr)$. This means we are interested in the covariance computed over an infinite run of SGD optimization, rather than the covariance at a fixed update step $k$ computed over multiple runs of SGD (see ˜A).
To better understand the behavior of the weight variances, we further examine the covariance matrix of the velocities, $\boldsymbol{\mathbf{\Sigma_{v}}}\coloneqq\textrm{cov}\bigl(\vec{v}_{k},\vec{v}_{k}\bigr)$. We then explore the ratio of the weight variance to the velocity variance in any given direction, which we denote as $\tau_{i}$. Under general assumptions, this ratio equates to the velocity correlation time of the corresponding direction (see Section˜4.2).
@@ 3 Related Work
3.1 Hessian and gradient sample covariance
The assumption of a strong alignment between the gradient sample covariance $\boldsymbol{\mathbf{C_{0}}}$ and the Hessian matrix of the training loss function $\boldsymbol{\mathbf{H}}$ is widely used in the literature [15, 22, 30]. Various theoretical arguments suggest that when a neural network’s output closely matches the example labels, these two matrices should be similar [15, 20, 21, 23, 34]. However, as highlighted by Thomas et al. [31], even slight deviations between network predictions and labels can theoretically disrupt this relationship. Nevertheless, empirical observations often reveal a strong alignment between the gradient sample covariance and the Hessian matrix near a minimum.
Zhang et al. [22] investigated this assumption numerically. They analyzed both matrices in a specific basis that exhibits varying curvature across different directions. Their findings showed a close correspondence between curvature and gradient variance along a given direction in a convolutional image recognition network, and a reasonably good relationship in a transformer model. Similarly, Thomas et al. [31] provided both theoretical arguments and empirical evidence across different image recognition architectures. While they did not find an exact match between the two matrices, they observed a proportionality indicated by a high cosine similarity. Xie et al. [17] also explored this relationship using an image recognition network. In the eigenspace of the Hessian matrix, they plotted entries within a specific interval against corresponding entries from the gradient sample covariance and observed a close match.
Ziyin et al. [25] emphasized the importance of considering the conditions under which the approximation of proportionality between the two matrices is valid, noting that the approximation may be questionable otherwise. However, the conditions they specify – particularly being near a minimum with low loss – are consistent with our empirical setup (see Section˜5.1 and Section˜G.1).
@@ 3.2 Limiting Dynamics and Weight Fluctuations
Several studies have examined the limiting dynamics of SGD, often modeling it as a stochastic differential equation (SDE). Researchers such as Mandt et al. [29] and Jastrzębski et al. [15] commonly approximate the loss near a minimum as a quadratic function, representing the SDE as a multivariate Ornstein-Uhlenbeck (OU) process. This process suggests a stationary weight distribution with Gaussian fluctuations. Jastrzębski et al. [15] further assume that the gradient covariance is proportional to the Hessian, observing under these conditions that the weight fluctuations are isotropic. Kunin et al. [28], who also incorporate momentum into their analysis, predict and empirically verify isotropic weight fluctuations. Chaudhari & Soatto [35] investigate the SDE without assuming a quadratic loss or equilibrium, gaining insights via the Fokker-Planck equation.
Alternatively, some studies derive relationships from a stationarity assumption rather than a continuous-time approximation [27, 30, 25]. Yaida [27] assumes that the weight trajectory follows a stationary distribution and derives general fluctuation-dissipation relations from this premise. Liu et al. [30] go further by assuming a quadratic loss function, enabling them to derive exact relations for the weight variance of SGD with momentum. If the gradient covariance is additionally assumed to be proportional to the Hessian, their results also predict that the weight variance is approximately isotropic, except in directions where the product of the learning rate and the Hessian eigenvalue is significantly large.
These theoretically computed weight variances are explicitly applied in various contexts, such as calculating the escape rate from a minimum or assessing the approximation error in SGD – which captures the additional training error attributed to noise [30].
Feng & Tu [32] present a phenomenological theory based on their empirical findings, which, unlike Kunin et al. [28], also accounts for flat directions. They describe a general inverse variance-flatness relationship by analyzing the weight trajectories of different image recognition networks via principal component analysis. They discovered a power law relationship between the curvature of the loss and the weight variance $\sigma_{\theta,i}^{2}$ in any given direction, where higher curvature corresponds to higher variance. They also observed that both the velocity variance $\sigma_{v,i}^{2}$ and the correlation time $\tau_{i}$ are larger for higher curvatures.
In our approach, we avoid the continuous-time approximation and instead base our results on the assumption that the weights adhere to a stationary distribution near a quadratic minimum.